A box rests on a frozen pond, which serves as a frictionless horizontal surface. If a fisherman applies a horizontal force with magnitude 48.0 N to the box and produces an acceleration of magnitude 3.00 m/s2, what is the mass of the box?


1
Expert's answer
2021-12-13T16:54:59-0500

Solution:


If the fisherman applies a horizontal force of magnitude 48 N, the produced acceleration on the box measures to be = 3 m/s2

thus from Newton's second law, we write

Net force "= mass \times acceleration"

"F = m \times a"

mass of the box "= \frac{Force }{ acceleration}""= \frac{48 }{ 3}" = 16 kg


Answer: mass of the box is 16 kg